# Beam Enumeration: Probabilistic Explainability For Sample Efficient Self-conditioned Molecular Design

### v1

@@ 5.1 Illustrative Experiment
Extracted Substructures are Meaningful. The illustrative experiment aims to optimize the following multi-parameter optimization (MPO) objective: maximize topological polar surface area (tPSA), molecular weight (MW) < 350 Da, and number of rings $\geq$ 2. This specific MPO was chosen because it is plausible to predict what structural features would be necessary to optimize the objective: rings saturated with heteroatoms. Due to constraining the molecular weight, the model cannot just learn to generate large molecules that would, on average, possess a higher tPSA. Augmented MemoryGuo and Schwaller (2023) was used to optimize the MPO objective. The reward trajectory tends towards 1, indicating the model gradually learns to satisfy the target objective, as desired (Fig. 2) Next, we investigate the top $k$ and $N$ beam steps parameters for Beam Enumeration and show that while the majority of sub-sequences do not possess valid substructures, a meaningful signal can still be extracted (Appendix C). We hypothesize that the optimal parameters are using a low top $k$ as we are interested in the most probable sub-sequences and large $N$ beam steps, which would enable extracting larger (and potentially more meaningful) substructures. Fig. 2 shows the top four substructures from Beam Enumeration at varying epochs. As hypothesized, the substructures are informative when considering the MPO objective: the most frequent substructures gradually become rings saturated with heteroatoms, which possess a high tPSA.
@@
Self-conditioned Generation Improves Sample Efficiency. Thus far, the results only show that Beam Enumeration can extract meaningful molecular substructures. To enable self-conditioned generation, a criterion is required to decide when to execute Beam Enumeration. We consider when extracted substructures would be meaningful and propose to execute Beam Enumeration when the reward improves for Patience number of successive epochs (to mitigate sampling stochasticity). We combine Beam Enumeration with Augmented MemoryGuo and Schwaller (2023) and perform an exhaustive hyperparameter grid search (with replicates) using Yield and Oracle Burden as the performance metrics (Appendix A). The results elucidate the behavior of Beam Enumeration with three key observations: firstly, Structure extraction is much more permissive compared to Scaffold and often leads to small functional groups, e.g., carbonyl, being the most frequent substructures which diminish the sample efficiency benefits (Appendix C). Secondly, enforcing larger substructures to be extracted (Structure Minimum Size) improves performance across all hyperparameter combinations. This reinforces that extracted substructures are meaningful as larger substructures heavily bias molecular generation during self-conditioning. If they were not meaningful, sample efficiency would not improve (and would likely be detrimental). Thirdly, Structure extraction while enforcing a higher Structure Minimum Size prevents small functional group extraction which significantly enhances performance.
@@
We perform five experiments (N=100 replicates each) based on the optimal hyperparameters identified from the grid search: Augmented MemoryGuo and Schwaller (2023) (baseline) and Augmented Memory with Beam Enumeration (Scaffold and Structure with and without Structure Minimum Size = 15). Table 1 shows that Beam Enumeration drastically improves the Yield and Oracle Burden compared to the baseline at both the > 0.7 and > 0.8 reward thresholds, particularly when Structure Minimum Size = 15 is enforced. We highlight that the improved sample efficiency is especially significant as baseline Augmented Memory could not find 100 molecules > 0.8 reward in 69/100 replicates. The results are further compared using Welch’s t-test, and all p-values are significant at the 95% confidence level (Appendix C).
@@ 5.2 Drug Discovery Case Studies
Figure 3: Three drug discovery case studies showing the top generated molecule (triplicate experiments) using Augmented MemoryGuo and Schwaller (2023) with Beam Enumeration Structure Minimum Structure Size = 15 and the reference ligand. Extracted substructures from Beam Enumeration are highlighted. The multi-parameter optimization objective is: Minimize Vina score, maximize QED, and molecular weight < 500 Da. a. Dopamine type 2 receptorWang et al. (2018). b. MK2 kinaseArgiriadi et al. (2010). c. AcetylcholinesteraseKryger et al. (1999).
@@
The positive results from the illustrative experiment suggest that Beam Enumeration can be applied to real-world drug discovery case studies to design inhibitors against DRD2 which is implicated in neurodegenerative diseasesWang et al. (2018), MK2 kinase which is involved in pro-inflammatory responsesArgiriadi et al. (2010), and AChE which is a target of interest against Alzheimer’s diseaseKryger et al. (1999). Following Guo et al.Guo et al. (2021a); Guo and Schwaller (2023), we formulate the following MPO objective: minimize the AutoDock VinaTrott and Olson (2010) docking score as a proxy for binding affinity, maximize the Quantitative Estimate of Druglikeness (QED)Bickerton et al. (2012) score, and constrain MW < 500 Da. The QED and MW objectives prevent the generative model from exploiting the weaknesses of docking algorithms to give inflated docking scores to large, lipophilic molecules, which can be promiscuous bindersArnott and Planey (2012). Moreover, an oracle budget of 5,000 Vina calls was enforced which is almost half the budget of the original Augmented MemoryGuo and Schwaller (2023) work (9,600).
@@
Since the observations made from the illustrative experiment hyperparameter grid search may not be generalizable to docking tasks, we perform an additional hyperparameter grid search (with replicates). The results (Appendix D) show that the optimal hyperparameters from the illustrative experiment are also the optimal hyperparameters across all three drug discovery case studies. We designate these the default hyperparameters and demonstrate the applicability of Beam Enumeration to both Augmented MemoryGuo and Schwaller (2023) and REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a) which is the second most (behind Augmented Memory) sample efficient model in the PMOGao et al. (2022) benchmark.
@@
Qualitative Inspection: Why Were These Molecules Generated? We first show that Augmented Memory with Beam Enumeration generates molecules that satisfy the MPO objective (Fig. 3). We emphasize that results were not cherry-picked and the three generated examples shown are the top 1 (by reward) across triplicate experiments. All molecules possess better Vina scores and higher QED than the reference molecules, as desired. Fig. 3 shows the highlighted substructures extracted using Structure extraction with Structure Minimum Size = 15 with three key observations: firstly, "uncommon" molecular substructures may be extracted such as the bridged cycle against DRD2. The exact substructure extracted was an amide bond with a long carbon chain which implicitly enforces the bridged cycle, and the Vina pose shows that it fits in the binding cavity with no clashes, despite being a bulky group. Secondly, bicylic or double-ring systems are often extracted (for all case studies), forming central scaffolds of the full molecule. Thirdly, scaffolds with branch points, i.e., a central ring with single carbon bond extensions, are often extracted (for all case studies). These substructures are particularly interesting as they heavily bias what can be generated in the remaining portion of the full molecule. An exemplary example of this is in the first generated molecule against MK2, where the branch points are effectively a part of two other ring systems (Fig. 3). Beam Enumeration can provide insights into the tolerability and suitability of certain substructures in the context of the full molecules (see Appendix D for more examples of extracted substructures). Finally, we posit that the extreme bias of Structure extraction is the reason why it can be more performant than Scaffold. Overall, the extracted substructures are meaningful and act both as a source of generative explainability and can self-direct the generative model into specific regions of chemical space with high reward.
@@
Quantitative Analysis: Sample Efficiency. Next, we reinforce results from previous work showing that Augmented MemoryGuo and Schwaller (2023) is significantly more sample efficient than REINVENTOlivecrona et al. (2017); Blaschke et al. (2020b) (Table 2). Notably, the Yield of Augmented Memory is much greater than REINVENT at both the > 0.7 and > 0.8 reward thresholds, indicating that more high reward molecules are generated. Moreover, Augmented Memory has a lower Oracle Burden than REINVENT in all cases, except for Oracle Burden>0.8 (1) for DRD2 and AChE where there is essentially no difference. The reason for this is because molecules with > 0.8 reward were already generated at epoch 1, indicating the pre-trained model (trained on ChEMBLGaulton et al. (2012a)) is a good Prior for these case studies. By contrast, the MK2 case study is considerably more challenging as extremely few > 0.8 reward molecules are generated under a 5,000 oracle calls budget. Augmented Memory significantly outperforms REINVENT as the latter could not find 10 molecules with reward > 0.8 (Table 2).
@@
Subsequently, we demonstrate that Beam Enumeration can be applied out-of-the-box on top of Augmented Memory and REINVENT. Firstly, the addition of Beam Enumeration improves the sample efficiency of both base algorithms, as evidenced by the Yield and Oracle Burden metrics in Table 2 with a small trade-off in diversity (Appendix D). However, the benefits are more pronounced in Augmented Memory as observed by the Yield>0.8 improving by > 4x in all cases (MK2 improves by 29x) and the Oracle Burden >0.8 (10 and 100) over halved in most cases. Notably, for MK2 Oracle Burden >0.8 (100), baseline Augmented Memory could not accomplish the task while Beam Enumeration is successful in almost under 2,000 oracle calls (Table 2). These findings are in agreement with the original Augmented MemoryGuo and Schwaller (2023) work in that the algorithm is much more data efficient and capitalizes on learning from high reward molecules via experience replayLin . Furthermore, we verify that a large number of unique scaffolds are generated despite the Beam Enumeration bias (Appendix D), demonstrating that the combined algorithm with Augmented Memory achieves both exploration and exploitation. Overall, the results show that Beam Enumeration is task-agnostic and can be applied on top of existing algorithms, including Augmented Memory and REINVENT, improving their sample efficiency. The combined algorithm generates more high reward molecules and faster, even in challenging (MK2) scenarios under a limited oracle budget. Furthermore, in reference to all the Oracle Burden metrics (Table 2), Augmented Memory with Beam Enumeration can identify a small set of excellent (high reward) candidate molecules in under 2,000 oracle calls and in some cases, even under 1,000 oracle calls.
@@
Oracle Burden>0.8 (100) (↓) DRD2 1120 $\pm$ 194 1056 $\pm$ 146 2558 $\pm$ $*{{}^{*}}$ 1928 $\pm$ 117 2109 $\pm$ 1090 4595 $\pm$ ${{}^{**}}$
MK2 2189 $\pm$ 181 2676 $\pm$ 403 Failed 3208 $\pm$ ${{}^{**}}$ Failed Failed
@@ 6 Conclusion
In this work, we propose Beam Enumeration to exhaustively enumerate sub-sequences from a language-based molecular generative model based on the top $k$ most probable tokens and for $N$ beam steps. We show that molecular substructures can be extracted from these sub-sequences, providing a source of generative explainability. Next, we show that the extracted molecular substructures can be used to self-condition the generative model to only send generated molecules possessing these substructures for oracle feedback (discarding the rest). We show that Beam Enumeration can be coupled with existing RL-based algorithms including Augmented MemoryGuo and Schwaller (2023) and REINVENTOlivecrona et al. (2017); Blaschke et al. (2020b) to improve their sample efficiency. Moreover, enforcing the extraction of larger substructures improves performance across all hyperparameter combinations. We believe this is a particularly interesting observation as it demonstrates the model’s remarkable robustness and tolerability to extreme bias. Subsequently, in three drug discovery case studies to design molecules that dock well, the addition of Beam Enumeration to Augmented Memory and REINVENT drastically improves sample efficiency as assessed by the Yield (number of unique molecules generated above a reward threshold) and Oracle Burden (number of oracle calls required for the model to generate $N$ unique molecules above a reward threshold) with a small trade-off in diversity (which is expected). The extracted substructures themselves provide valuable structural insights, often enforcing the generation of specific cyclic systems and scaffolds with branch points which impose an overall molecular geometry that complements the protein binding cavity, thus serving as a source of generative explainability. Beam Enumeration is the first proposed method to jointly address explainability and sample efficiency in molecular generative models. The improvements in the latter will enable more expensive high-fidelity oracles to be explicitly optimized. We note, however, that sparse reward environmentsKorshunova et al. (2022) remain a difficult optimization task. Finally, Beam Enumeration is a task-agnostic method and can be combined with recent work integrating active learning with molecular generation to further improve sample efficiencyDodds et al. (2023); Kyro et al. (2023). If the benefits can be synergistic, we may approach sufficient sample efficiency to directly optimize expensive state-of-the-art (in predictive accuracy) physics-based oracles such as MD simulationsWang et al. (2015); Moore et al. (2023). Excitingly, this would in turn enhance explainability as high-fidelity oracles are inherently more informative.
Data and code availability
@@
All code and prepared configuration files to reproduce all experiments in the main text is provided in the following repository: https://github.com/schwallergroup/augmented_memory.
Acknowledgement
This publication was created as part of NCCR Catalysis (grant number 180544), a National Centre of Competence in Research funded by the Swiss National Science Foundation.
Appendix
@@ Acknowledgement
Appendix A Beam Enumeration

### v2

@@ 5.1 Illustrative Experiment
Extracted Substructures are Meaningful. The illustrative experiment aims to optimize the following multi-parameter optimization (MPO) objective: maximize topological polar surface area (tPSA), molecular weight (MW) <350 Da, and number of rings $\geq$ 2. This specific MPO was chosen because it is plausible to predict what structural features would be necessary to optimize the objective: rings saturated with heteroatoms. Due to constraining the molecular weight, the model cannot just learn to generate large molecules that would, on average, possess a higher tPSA. Augmented Memory (Guo & Schwaller (2023) was used to optimize the MPO objective. The reward trajectory tends towards 1, indicating the model gradually learns to satisfy the target objective, as desired (Fig. 2) Next, we investigate the top $k$ and $N$ beam steps parameters for Beam Enumeration and show that while the majority of sub-sequences do not possess valid substructures, a meaningful signal can still be extracted (Appendix D). We hypothesize that the optimal parameters are using a low top $k$ as we are interested in the most probable sub-sequences and large $N$ beam steps, which would enable extracting larger (and potentially more meaningful) substructures. Fig. 2 shows the top four substructures from Beam Enumeration at varying epochs. As hypothesized, the substructures are informative when considering the MPO objective: the most frequent substructures gradually become rings saturated with heteroatoms, which possess a high tPSA.
@@
Self-conditioned Generation Improves Sample Efficiency. Thus far, the results only show that Beam Enumeration can extract meaningful molecular substructures. To enable self-conditioned generation, a criterion is required to decide when to execute Beam Enumeration. We consider when extracted substructures would be meaningful and propose to execute Beam Enumeration when the reward improves for Patience number of successive epochs (to mitigate sampling stochasticity). We combine Beam Enumeration with Augmented Memory (Guo & Schwaller (2023) and perform an exhaustive hyperparameter grid search (with replicates) using Yield and Oracle Burden as the performance metrics (Appendix B). The results elucidate the behavior of Beam Enumeration with three key observations: firstly, Structure extraction is much more permissive compared to Scaffold and often leads to small functional groups, e.g., carbonyl, being the most frequent substructures which diminish the sample efficiency benefits (Appendix D). Secondly, enforcing larger substructures to be extracted (Structure Minimum Size) improves performance across all hyperparameter combinations. This reinforces that extracted substructures are meaningful as larger substructures heavily bias molecular generation during self-conditioning. If they were not meaningful, sample efficiency would not improve (and would likely be detrimental). Thirdly, Structure extraction while enforcing a higher Structure Minimum Size prevents small functional group extraction which significantly enhances performance.
@@
We perform five experiments (N=100 replicates each) based on the optimal hyperparameters identified from the grid search: Augmented Memory (Guo & Schwaller (2023) (baseline) and Augmented Memory with Beam Enumeration (Scaffold and Structure with and without Structure Minimum Size = 15). Table 1 shows that Beam Enumeration drastically improves the Yield and Oracle Burden compared to the baseline at both the >0.7 and >0.8 reward thresholds, particularly when Structure Minimum Size = 15 is enforced. We highlight that the improved sample efficiency is especially significant as baseline Augmented Memory could not find 100 molecules >0.8 reward in 69/100 replicates. The results are further compared using Welch’s t-test, and all p-values are significant at the 95% confidence level (Appendix D).
@@ 5.2 Drug Discovery Case Studies
Figure 3: Three drug discovery case studies showing the top generated molecule (triplicate experiments) using Augmented Memory (Guo & Schwaller (2023) with Beam Enumeration Structure Minimum Structure Size = 15 and the reference ligand. Extracted substructures from Beam Enumeration are highlighted. The multi-parameter optimization objective is: Minimize Vina score, maximize QED, and molecular weight <500 Da. The values, with the Synthetic Accessibility (SA) score (Ertl & Schuffenhauer (2009) are annotated. a. Dopamine type 2 receptor (Wang et al. (2018). b. MK2 kinase (Argiriadi et al. (2010). c. Acetylcholinesterase (Kryger et al. (1999).
@@
The positive results from the illustrative experiment suggest that Beam Enumeration can be applied to real-world drug discovery case studies to design inhibitors against DRD2 which is implicated in neurodegenerative diseases (Wang et al. (2018), MK2 kinase which is involved in pro-inflammatory responses (Argiriadi et al. (2010), and AChE which is a target of interest against Alzheimer’s disease (Kryger et al. (1999). Following Guo et al. (Guo et al. (2021b); Guo & Schwaller (2023), we formulate the following MPO objective: minimize the AutoDock Vina (Trott & Olson (2010) docking score as a proxy for binding affinity, maximize the Quantitative Estimate of Druglikeness (QED) (Bickerton et al. (2012) score, and constrain MW <500 Da. The QED and MW objectives prevent the generative model from exploiting the weaknesses of docking algorithms to give inflated docking scores to large, lipophilic molecules, which can be promiscuous binders (Arnott & Planey (2012). Moreover, an oracle budget of 5,000 Vina calls was enforced which is almost half the budget of the original Augmented Memory (Guo & Schwaller (2023) work (9,600).
@@
Since the observations made from the illustrative experiment hyperparameter grid search may not be generalizable to docking tasks, we perform an additional hyperparameter grid search (with replicates). The results (Appendix E) show that the optimal hyperparameters from the illustrative experiment are also the optimal hyperparameters across all three drug discovery case studies. We designate these the default hyperparameters and demonstrate the applicability of Beam Enumeration to both Augmented Memory (Guo & Schwaller (2023) and REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020) which is the second most (behind Augmented Memory) sample efficient model in the PMO (Gao et al. (2022) benchmark.
@@
Qualitative Inspection: Why Were These Molecules Generated? We first show that Augmented Memory with Beam Enumeration generates molecules that satisfy the MPO objective (Fig. 3). We emphasize that results were not cherry-picked and the three generated examples shown are the top 1 (by reward) across triplicate experiments. All molecules possess better Vina scores and higher QED than the reference molecules, as desired. Fig. 3 shows the highlighted substructures extracted using Structure extraction with Structure Minimum Size = 15 with three key observations: firstly, ”uncommon” molecular substructures may be extracted such as the bridged cycle against DRD2. The exact substructure extracted was an amide bond with a long carbon chain which implicitly enforces the bridged cycle, and the Vina pose shows that it fits in the binding cavity with no clashes, despite being a bulky group. Secondly, bicylic or double-ring systems are often extracted (for all case studies), forming central scaffolds of the full molecule. Thirdly, scaffolds with branch points, i.e., a central ring with single carbon bond extensions, are often extracted (for all case studies). These substructures are particularly interesting as they heavily bias what can be generated in the remaining portion of the full molecule. An exemplary example of this is in the first generated molecule against MK2, where the branch points are effectively a part of two other ring systems (Fig. 3). Beam Enumeration can provide insights into the tolerability and suitability of certain substructures in the context of the full molecules (see Appendix E for more examples of extracted substructures). Finally, we posit that the extreme bias of Structure extraction is the reason why it can be more performant than Scaffold. Overall, the extracted substructures are meaningful and act both as a source of generative explainability and can self-direct the generative model into specific regions of chemical space with high reward.
@@
Quantitative Analysis: Sample Efficiency. Next, we reinforce results from previous work showing that Augmented Memory (Guo & Schwaller (2023) is significantly more sample efficient than REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020) (Table 2). Notably, the Yield of Augmented Memory is much greater than REINVENT at both the >0.7 and >0.8 reward thresholds, indicating that more high reward molecules are generated. Moreover, Augmented Memory has a lower Oracle Burden than REINVENT in all cases, except for Oracle Burden>0.8 (1) for DRD2 and AChE where there is essentially no difference. The reason for this is because molecules with >0.8 reward were already generated at epoch 1, indicating the pre-trained model (trained on ChEMBL (Gaulton et al. (2012a)) is a good Prior for these case studies. By contrast, the MK2 case study is considerably more challenging as extremely few >0.8 reward molecules are generated under a 5,000 oracle calls budget. Augmented Memory significantly outperforms REINVENT as the latter could not find 10 molecules with reward >0.8 (Table 2).
@@
Subsequently, we demonstrate that Beam Enumeration can be applied out-of-the-box on top of Augmented Memory and REINVENT. Firstly, the addition of Beam Enumeration improves the sample efficiency of both base algorithms, as evidenced by the Yield and Oracle Burden metrics in Table 2 with a small trade-off in diversity (Appendix E). However, the benefits are more pronounced in Augmented Memory as observed by the Yield>0.8 improving by >4x in all cases (MK2 improves by 29x) and the Oracle Burden >0.8 (10 and 100) over halved in most cases. Notably, for MK2 Oracle Burden >0.8 (100), baseline Augmented Memory could not accomplish the task while Beam Enumeration is successful in almost under 2,000 oracle calls (Table 2). These findings are in agreement with the original Augmented Memory (Guo & Schwaller (2023) work in that the algorithm is much more data efficient and capitalizes on learning from high reward molecules via experience replay (Lin (1992). Beam Enumeration decreases the diversity of the generated set (as measured by IntDiv1Polykovskiy et al. (2020b), but finds considerably more unique scaffolds above the 0.8 reward threshold (up to 19x) (Appendix E). With many unique scaffolds built around (often) central substructures, the generated set could conceivably provide insights into structure-activity relationships. These results demonstrate that the combined algorithm achieves both exploration and exploitation. Overall, the results show that Beam Enumeration is task-agnostic and can be applied on top of existing algorithms to improve their sample efficiency. The combined algorithms generate more high reward molecules and faster, even in challenging (MK2) scenarios under a limited oracle budget. Furthermore, in reference to all the Oracle Burden metrics (Table 2), Augmented Memory with Beam Enumeration can identify a small set of excellent (high reward) candidate molecules in under 2,000 oracle calls and in some cases, even under 1,000 oracle calls.
@@
Oracle Burden>0.8 (100) (↓) DRD2 1120 $\pm$ 194 1056 $\pm$ 146 2558 $\pm$ $*{{}^{*}}$ 1928 $\pm$ $*{{}^{*}}$ 2109 $\pm$ 1090 4595 $\pm$ ${{}^{**}}$
MK2 2189 $\pm$ 181 2676 $\pm$ $*{{}^{*}}$ Failed 3208 $\pm$ ${{}^{**}}$ Failed Failed
@@ 6 Conclusion
In this work, we propose Beam Enumeration to exhaustively enumerate sub-sequences from language-based molecular generative models and show that substructures can be extracted, providing a source of generative explainability. Next, we show that the extracted molecular substructures can be used to self-condition the generative model to only perform oracle evaluation for molecules possessing these substructures (discarding the rest). We show that Beam Enumeration can be coupled with existing RL-based algorithms including Augmented Memory (Guo & Schwaller (2023) and REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020) to improve their sample efficiency. Moreover, enforcing the extraction of larger substructures improves performance across all hyperparameter combinations. We believe this is a particularly interesting observation as it demonstrates the model’s remarkable robustness and tolerability to extreme bias. Subsequently, in three drug discovery case studies to design molecules that dock well, the addition of Beam Enumeration to Augmented Memory and REINVENT substantially improves sample efficiency as assessed by the Yield (number of unique molecules generated above a reward threshold) and Oracle Burden (number of oracle calls required for the model to generate $N$ unique molecules above a reward threshold) with a small trade-off in diversity (which is expected). The extracted substructures themselves provide valuable structural insights, often enforcing the generation of specific cyclic systems and scaffolds with branch points which impose an overall molecular geometry that complements the protein binding cavity. Beam Enumeration shows that improvements to explainability and sample efficiency for molecular design can be made synergistic. The improvements in the latter will enable more expensive high-fidelity oracles to be explicitly optimized. We note, however, that sparse reward environments (Korshunova et al. (2022) remain a difficult optimization task. Finally, Beam Enumeration is a task-agnostic method and can be combined with recent work integrating active learning with molecular generation to further improve sample efficiency (Dodds et al. (2024); Kyro et al. (2023). If the benefits can be synergistic, we may approach sufficient sample efficiency to directly optimize expensive state-of-the-art (in predictive accuracy) physics-based oracles such as MD simulations (Wang et al. (2015); Moore et al. (2023). Excitingly, this would in turn enhance explainability as high-fidelity oracles are inherently more informative.
7 Reproducibility Statement
@@
The code is provided in the GitHub link in the Abstract and also provided here: https://github.com/schwallergroup/augmented_memory. In the repository, there are prepared configuration files that can be directly run to reproduce all experiments in this work.
Appendix A appendix
@@
Appendix B Beam Enumeration